OpenStudy (emily_98):

John has 48 square centimeter tiles he wants to use to create a mosaic. He wants the mosaic to be rectangular with a length that is 2 centimeters longer than the width

OpenStudy (emily_98):

Which equation could John solve to find w, the greatest width in centimeters he can use for the mosaic? w(w – 2) = 48 w(w + 2) = 48 2w(w – 2) = 48 2w(w + 2) = 48

OpenStudy (emily_98):

@Astrophysics